A charge of 6x10-7C is transferred from infinity to point A. If the work done to do this is 1.2x10-5J what is the potential point A


1
Expert's answer
2022-04-05T19:18:58-0400

The potential at point A:

"V=\frac{W}{q}=\rm\frac{1.2*10^{-5}\: J}{6*10^{-7}\: C}=20\: V"